Mcnamee Corporation's activity-based costing system has three activity cost pools--Fabricating, Setting Up, and Other. Costs in the Fabricating cost pool are assigned to products based on machine-hours (MHs) and costs in the Setting Up cost pool are assigned to products based on the number of batches. Costs in the Other cost pool are not assigned to products. The activity rate for Fabricating is $0.18 per MH; the activity rate for Setting Up is $3.34 per batch. The following table shows the machine-hours and number of batches associated with each of the company's two products:

MHs Batches
Product O0 2,600 240
Product O1 17,000 1,760
Total 19,600 2,000

Required:

Determine the amount of overhead cost that would be assigned to each product using activity-based costing.

Amount of Overhead Cost
Product O0
Product O1
Answer #1

Answer:

Product 00 Product 01
Fabricating 468 3060
Setting up 802 5878
Overhead cost assigned 1270 8938
Workings:
Product 00 Product 01
Fabricating =2600*0.18 =17000*0.18
Setting up =240*3.34 =1760*3.34